OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -one halfx + 5. y = -2x + 2 y = 2x - 14 y = one halfx - 8 y = -one halfx - 4

OpenStudy (anonymous):

y= -1/2x -4 because for them to be parallel, the slopes have to be the same. you would take the points you were given and figure x=4 and y=-6 because of the coordinate point. Then you would plug -6 in as your y in the equation y=mx+b M=-1/2 y=-6 x=4 Then you just do simple algebra and you would then get you value for b. and plug it back in y=-1/2x+? And there is your answer. Hope I helped!

OpenStudy (anonymous):

Thank you I thought it was that I just wasn't sure.

OpenStudy (anonymous):

No problem. (: